OpenStudy (anonymous):

3-(a-4/a+4)=a^2-8/a+4

myininaya (myininaya):

multiply both sides by a+4

OpenStudy (anonymous):

when I did that I came up with 3a+16=a^2-8

OpenStudy (anonymous):

I am currently stuck at 0=a^2-3a-24

myininaya (myininaya):

3(a+4)-(a-4)=a^2-8 3a+12-a+4=a^2-8 2a+16=a^2-8 0=a^2-2a-8-16 0=a^2-2a-24 is what i get

OpenStudy (anonymous):

may I just ask how you got the 2a, I knew that I had messed up somewhere, but I kept getting 3a and 24

myininaya (myininaya):

3a-a

OpenStudy (anonymous):

but where does that other a come from?

myininaya (myininaya):

3a-a=2a then i subtracted 2a on both sides

myininaya (myininaya):

i multiplied both sides by a+4 3(a+4)-(a-4)=a^2-8

OpenStudy (anonymous):

Thank you!! I couldnt figure out where I went wrong !!

OpenStudy (anonymous):

how do I break it down from the point of 0=a^2-2a-24? I divided by 2 and got 12, but the final answer is supposed to be 6, could you tell me where I went wrong?

myininaya (myininaya):

to factor a^2-2a-24 find two factors of -24 that have product -24 and have sum -2

OpenStudy (anonymous):

6 and 4?

myininaya (myininaya):

-6 and 4 -6(4)=-24 -6+4=-2 so we have (x-6)(x+4)=0

myininaya (myininaya):

remember we don't want a=-4 since neither function on either side is defined at x=-4 only solution is a=6

OpenStudy (anonymous):

thank you again :]